OpenStudy (anonymous):

I have about 5 questions i need some major help on.

OpenStudy (anonymous):

What values for \[\theta (0\le \theta \le 2n)\] Satisfy the equation? \[2\sin \theta \cos \theta +\sqrt{3} \cos \theta = 0\]

OpenStudy (anonymous):

\[\frac{ \pi }{ 2 } \frac{ 4\pi }{ 3 }\ \frac{ 3\pi }{2}\frac{ 5\pi }{ 3 }\] \[\frac{ \pi }{ 2 }\frac{ 3\pi }{ 4 }\frac{ 3\pi }{ 4 }\frac{ 5\pi}{ 3 }\]

OpenStudy (anonymous):

\[\frac{ \pi }{ 2 }\frac{ 3\pi }{ 4 }\frac{ 3\pi }{ 2}\frac{ 5\pi }{ 4 }\]

OpenStudy (anonymous):

\[\frac{ \pi }{ 2 }\frac{ \pi}{ 4}\frac{ 3\pi}{ 2 }\frac{ 5\pi }{ 3 }\]

OpenStudy (sleepyjess):

@agent0smith

OpenStudy (cggurumanjunath):